OpenStudy (maddix):

Can Some one help me With This 90 divided 20 =

OpenStudy (mecharv):

\[\frac{ 90 }{ 20 }\] is this what you need help with?

OpenStudy (maddix):

yes

OpenStudy (mecharv):

\[\frac{ 90 }{ 20 }= \frac{ 9 }{ 2 }=4.5\]

OpenStudy (maddix):

Thanks

OpenStudy (mecharv):

Anytime :)

OpenStudy (tocuteforyou2):

can someone help me with history plz

OpenStudy (maddix):

yes

OpenStudy (tocuteforyou2):

should i post it here or no